キャリブレーション結果

2006 年 4 月 23 日

[figure]

• レートテーブルで測定 • 測定日は 2006/04/10 • 0 号機: ADXRS150 1 号機: CRS03

[figure]

図 1 0 号機 Mode1 図 2 0 号機 Mode2 図 3 0 号機 Mode3 図 4 1 号機 Mode1 図 5 1 号機 Mode1 図 6 1 号機 Mode3

4